OpenStudy (pflori1234_pop):

Anyone really good with inequalities... i have the answers to the questions i just need the steps

OpenStudy (pflori1234_pop):

7- 2(x-4) < 5(1 -2x)

OpenStudy (pflori1234_pop):

the answer is: x < -5/4

alones (alones):

You need the steps again?

OpenStudy (pflori1234_pop):

yea

OpenStudy (pflori1234_pop):

cuz I'm studying for my test tomorrow and my teacher gave us a study guide with the questions and the answers and some of them i can't figure it out

OpenStudy (mykal):

ok first you should simplify

OpenStudy (mykal):

ultimately youre gonna wanna combine like terms

OpenStudy (mykal):

so first simplify both sides then combine like terms

jimthompson5910 (jim_thompson5910):

Algebraic Steps Step 1) 7- 2(x-4) < 5(1 -2x) Step 2) 7-2x+8 < 5 - 10x Step 3) -2x+15 < 5 - 10x Step 4) -2x+15+10x < 5 - 10x+10x Step 5) 8x+15 < 5 Step 6) 8x+15-15 < 5-15 Step 7) 8x < -10 Step 8) 8x/8 < -10/8 Step 9) x < -5/4 Explanation of each step (shown above) Step 1) Original inequality; no explanation needed really Step 2) Distribute. Make sure to distribute the outer term to EVERY term inside Step 3) Combine like terms (left side: 7+8 = 15) Step 4) Add 10x to both sides Step 5) Combine like terms (-2x+10x=8x and -10x+10x=0x) Step 6) Subtract 15 from both sides Step 7) Combine like terms (15-15 = 0 and 5-15 = -10) Step 8) Divide both sides by 8 to completely isolate x Step 9) Reduce all fractions as much as possible So when we solve `7- 2(x-4) < 5(1 -2x)` for x, we get `x < -5/4`